# E2E-MFD: Towards End-to-End Synchronous Multimodal Fusion Detection

### v2

@@ Abstract
Multimodal image fusion and object detection are crucial for autonomous driving. While current methods have advanced the fusion of texture details and semantic information, their complex training processes hinder broader applications. Addressing this challenge, we introduce E2E-MFD, a novel end-to-end algorithm for multimodal fusion detection. E2E-MFD streamlines the process, achieving high performance with a single training phase. It employs synchronous joint optimization across components to avoid suboptimal solutions tied to individual tasks. Furthermore, it implements a comprehensive optimization strategy in the gradient matrix for shared parameters, ensuring convergence to an optimal fusion detection configuration. Our extensive testing on multiple public datasets reveals E2E-MFD’s superior capabilities, showcasing not only visually appealing image fusion but also impressive detection outcomes, such as a 3.9% and 2.0% $\text{mAP}_{50}$ increase on horizontal object detection dataset M3FD and oriented object detection dataset DroneVehicle, respectively, compared to state-of-the-art approaches.
@@ 1 Introduction
Precise and reliable object parsing is critical in fields such as autonomous driving and remote sensing monitoring [1]. Relying solely on visible sensors can lead to inaccuracies in object recognition in challenging environments, like inclement weather conditions. Visible-infrared image fusion [2; 3; 4; 5] as a typical common multimodal fusion (MF) task addresses these challenges by leveraging complementary information from different modalities, leading to the rapid development of various multimodal image fusion techniques [6; 7; 8; 9; 10]. Techniques like CDDFuse [11] and DIDFuse [12] employ a two-step process where a MF network is trained initially, followed by training an object detection (OD) network with the results from the MF network to assess fusion effectiveness separately. Although deep neural networks have significantly enhanced the ability to learn representations across modalities, resulting in promising multimodal fusion outcomes, the focus has predominantly been on producing visually appealing images. This emphasis often overlooks the improvement of downstream high-level visual tasks, such as enhanced object parsing, which remains a substantial hurdle.
@@
Recent studies have ventured into designing joint learning methods that integrate fusion networks with high-level tasks such as object detection [13] and segmentation [14; 15]. The synergy between MF and OD in Multimodal Fusion Detection (MFD) methods has emerged as a vibrant area of research. This partnership allows MF to produce richer, more informative images, enhancing OD performance, while OD contributes valuable object semantic insights to MF, aiming to accurately locate and identify objects in a scene. Typically, MFD networks adopt a cascaded design where joint optimization techniques [16] use the OD network to guide the MF network toward creating images that facilitate easier object detection. Notably, Zhao et. al [17] introduced a joint learning method for multimodal fusion detection, incorporating meta-feature embedding from OD to improve fusion by generating semantic object features through meta-learning simulation. Despite these advancements, as highlighted in Figure 1, significant challenges persist: 1) Current optimization approaches rely on a multi-step, progressively joint method, compromising efficiency; 2) These methods overly focus on leveraging OD information for fusion enhancement, leading to difficulty in parameter balancing and susceptibility to local optima of individual tasks. Therefore, the quest for a unified feature set that simultaneously caters to each task remains formidable.
@@
In this paper, we introduce E2E-MFD, an end-to-end algorithm for multimodal fusion detection, designed to seamlessly blend detailed image fusion and object detection from coarse to fine levels. E2E-MFD facilitates the interaction of intrinsic features from both domains through synchronous joint optimization, allowing for a streamlined, one-stage process. To reconcile fine-grained details with semantic information, we propose the novel concept of an Object-Region-Pixel Phylogenetic Tree (ORPPT) coupled with a coarse-to-fine diffusion processing (CFDP) mechanism. This approach is inspired by the natural process of visual perception, tailored to meet the specific needs of MF and OD. Furthermore, we introduce a Gradient Matrix Task-Alignment (GMTA) technique to fine-tune the optimization of shared components, thereby minimizing the adverse impacts traditionally associated with inherent optimization challenges. This ensures an efficient convergence towards an optimal set of fusion detection weights, enhancing both the accuracy and efficacy of multimodal fusion detection tasks.
@@
Our contributions in this paper are highlighted as follows: (1) We present E2E-MFD, a pioneering approach to efficient synchronous joint learning, innovatively integrating image fusion and object detection into a single-stage, end-to-end framework. This methodology significantly enhances the outcomes of both tasks. (2) We introduce a novel Gradient Matrix Task-Alignment (GMTA) technique, designed to evaluate and quantify the impacts of the image fusion and object detection tasks. This aids in optimizing the training process’s stability and ensures convergence to an optimal configuration of fusion detection weights. (3) Through comprehensive experimentation on image fusion and object detection, we demonstrate the efficacy and robustness of our proposed method.
@@ 2.1 Multimodal Fusion Object Detection
Due to the powerful nonlinear fitting capabilities of deep neural networks, deep learning has made significant progress in low-level vision tasks, particularly in image fusion tasks [18; 19; 6; 20; 21; 22; 23]. Early efforts [8; 24; 12; 25; 26] tended to achieve excellent fusion results by adjusting network structures or loss functions, overlooking the fact that image fusion should aim to improve the performance of downstream application tasks. Fusion images with good quality metrics may be suitable for human visual perception but may not be conducive to practical application tasks [15]. Some research has acknowledged this issue, Yuan et al[27] utilized various aligned modalities for improved oriented object detection [28; 29; 30; 31] to address the challenge of cross-modal weakly misalignment in aerial visible-infrared images. Liu et al. [16] proposing a joint learning method, pioneered the exploration of the MF and OD combination methods. Then the optimization of the loss function for segmentation [15] and detection [13] has been validated to be effective in guiding the generation of fused images. They consider the downstream task network as an additional constraint to assist the MF network in generating fusion results with clearer objects. Zhao et al. [17] leveraged semantic information from OD features to aid MF and perform meta-feature embedding to generate meta-features from OD features, which are then used to guide the MF network in learning pixel-level semantic information. Liu et al. [14] proposed multi-interactive feature learning architecture for image fusion and segmentation by enhancing fine-grained mapping of all the vital information between two tasks, so that the modality or semantic features can be fully mutual-interactive. However, OD considers the semantic understanding of objects, while MF and segmentation primarily focus on the pixel-level relationship between image pairs. The optimization coupling of detection and fusion tasks becomes more challenging to investigate these complementary differences, from which both fusion and detection can benefit. It is worth mentioning that these visible-infrared multimodal fusion detection methods are usually designed in a cascaded structure with tedious training steps. Researchers lack the adoption of end-to-end architectures, which would enable one-step network inference to generate credible fused images and detection results through a set of network parameters.
@@ 2.2 Multi-task Learning
Multi-task learning (MTL) [32; 33] involves the simultaneous learning of multiple tasks through parameter sharing. Prior approaches involve manually crafting the architecture, wherein the bottom layers of a model are shared across tasks [34; 35]. Some approaches tailor the architecture based on task affinity [36], while others utilize techniques such as Neural Architecture Search [37; 34; 38] or routing networks [39] to autonomously discern sharing patterns and determine the architecture. An alternative method typically combines task-specific objectives into a weighted sum [40; 41; 42]. In addition, most approaches (e.g. [43; 44; 45; 46; 47]) aim to mitigate the effects of conflicting or dominating gradients. The approach of explicit gradient modulation [45; 46; 47; 48], has demonstrated superior performance in resolving conflicts between task gradients by substituting conflicting gradients with modified, non-conflicting gradients. Inspired by the above multimodal learning methods, we introduce a Gradient Matrix Task-Alignment method to align the orthogonal components contained in image fusion and object detection tasks thereby effectively eliminating the inherent optimization barrier that exists between two tasks.
@@ 3.1 Problem Formulation
MF-OD task concentrates on generating an image that benefits emphasizing objects with superior visual perception capability. The goal of OD is to find the location and identify the class of each object in an image which can naturally provide rich semantic information along with object location information. Therefore, the motivation of OD-aware MF is to construct a novel infrared and visible image fusion framework that can benefit from the semantic information and object location information contained in object detection. For this purpose, we suppose a pair of visible image $\boldsymbol{x}\in\mathbb{R}^{H\times W\times C_{x}}$ and infrared image $\boldsymbol{y}\in\mathbb{R}^{H\times W\times C_{y}}$. The optimization model is formulated as:
@@
where $\boldsymbol{t}$ represents the output of the different task network $\mathcal{N}$ with the learnable parameters $\boldsymbol{\theta}_{\mathrm{t}}$. $\mathcal{L}(\cdot)$ is a constraint term to optimize the network. Previous approaches solely design image fusion or object detection networks in a cascaded way, which can only achieve outstanding results for one task. To produce visually appealing fused images alongside accurate object detection results, we jointly integrate the two tasks into a unified goal synchronously which can be rewritten as:
@@ 3.2 Architecture
where $b_{l,m}(\boldsymbol{o_{l}})$ denotes the $m^{\mathrm{th}}$ region representation and $(i,j)$ denotes the spatial location. These region-level representations are further concatenated to form the observation $B_{l}(\boldsymbol{o_{l}})=\left[b_{l,1}(\boldsymbol{o_{l}}),b_{l,2}(\boldsymbol{o_{l}}),\ldots,b_{l,M_{l}}(\boldsymbol{o_{l}})\right]$ of branch $l$. These multi-grained attentions concentrate operation on the spatial location information and the extent of the regions which are similar to the task requirements. The $B_{1}$, $B_{2}$, …, $B_{L}$ are up-sampled to keep the consistent spatial size with the pixel-level fusion features $B_{0}$. Then, the region-level fusion features $B_{1},B_{2},...,B_{L}$ are assembled by addition operation followed by a Convolution with $1\times 1$ kernel + ReLU to reduce the channel numbers. Highlighted region features are extracted by a Convolution with $1\times 1$ kernel + Sigmoid and then injected into pixel-level features by multiplication and addition. Finally, five Convolutions with $3\times 3$ kernel + ReLU layers are constructed to reconstruct the fusion result $u$.
@@ 3.3 Loss Function
The total loss is combined with an image fusion loss function $\mathcal{L}_{\mathrm{f}}$ and object detection loss $\mathcal{L}_{\mathrm{d}}$. $\mathcal{L}_{\mathrm{f}}$ consists of three types of losses, i.e., structure loss $\mathcal{L}_{\mathrm{SSIM}}$, pixel loss $\mathcal{L}_{\text{pixel }}$ and gradient loss $\mathcal{L}_{\text{grad }}$. For one fused image, it should preserve overall structures and maintain a similar intensity distribution from source images. To this end, the structural similarity index (SSIM) is introduced in function:
@@
$\mathcal{L}_{\text{pixel}}^{\text{o}}=\left\|I_{m}\circ\left(\boldsymbol{u}-max(\boldsymbol{\omega_{\mathbf{1}}}\boldsymbol{x},\boldsymbol{\omega_{\mathbf{2}}}\boldsymbol{y})\right)\right\|_{1},\mathcal{L}_{\text{pixel}}^{\text{b}}=\left\|(1-I_{m})\circ\left(\boldsymbol{u}-mean(\boldsymbol{\omega_{\mathbf{1}}}\boldsymbol{x},\boldsymbol{\omega_{\mathbf{2}}}\boldsymbol{y})\right)\right\|_{1},$ (8)
where $\|\cdot\|_{1}$ stands for the $l_{1}$-norm. $\boldsymbol{\omega_{\mathbf{1}}}$ and $\boldsymbol{\omega_{\mathbf{2}}}$ are saliency weight maps calculated by $\boldsymbol{\omega}_{\mathbf{1}}=\boldsymbol{S}_{\boldsymbol{x}}(k)/\left[\boldsymbol{S}_{\boldsymbol{x}}(k)-\boldsymbol{S}_{\boldsymbol{y}}(k)\right],\boldsymbol{\omega}_{\mathbf{2}}=1-\boldsymbol{\omega}_{\mathbf{1}}$. Supposing that the saliency value of $\boldsymbol{x}$ at the $k^{\mathrm{th}}$ pixel can be obtained by $\boldsymbol{S}_{\boldsymbol{x}}(k)=\sum_{i=0}^{255}\boldsymbol{H}_{\boldsymbol{x}}(i)|\boldsymbol{x}(k)-i|$, where $\boldsymbol{x}(k)$ is the value of the $k$th pixel and $\boldsymbol{H}_{\boldsymbol{x}}(i)$ is the histogram of pixel value $i$. The operator $\circ$ denotes the elementwise multiplication, max($\cdot$) denotes the element-wise maximization, and mean($\cdot$) denotes the element-wise average operation. Therefore, the object-aware pixel loss $\mathcal{L}_{\text{pixel }}$ is defined as:
@@ 3.4 Gradient Matrix Task-Alignment
The MF and OD tasks have distinct optimization objectives. MF primarily emphasizes capturing the pixel-level relationship between image pairs, while OD incorporates object semantics within the broader context of diverse scenes. An inherent optimization barrier exists between these two tasks. We observe that the prevailing challenges in multi-task learning are arguably task dominance and conflicting gradients. We introduce a Gradient Matrix Task-Alignment (GMTA) by presenting the condition number to mitigate the undesired effects of the optimization barrier in task-shared parameters $\boldsymbol{\theta}^{\star}$ which are supposed to be balanced between the MF and OD tasks. The individual task gradients of MF and OD task are calculated by $\boldsymbol{g}_{u}=\nabla_{{\theta}^{\star}}\mathcal{L}_{u}$ and $\boldsymbol{g}_{d}=\nabla_{{\theta}^{\star}}\mathcal{L}_{d}$ in the training optimization process. The gradient matrix can be defined as $\boldsymbol{G}=\left\{\boldsymbol{g}_{u},\boldsymbol{g}_{d}\right\}$. In multi-task optimization, a cumulative gradient $\boldsymbol{g}=\boldsymbol{G}\boldsymbol{w}$ is a linear combination of task gradients and the stability of a linear system is measured by the condition number of its matrix in numerical analysis. Hence the stability of the gradient matrix is equal to the ratio of the maximum and minimum singular values of the corresponding matrix: $\kappa(\boldsymbol{G})=\frac{\sigma_{\max}}{\sigma_{\min}}$. Learning from the Aligned-MTL [38], The condition number is optimal ($\kappa(\boldsymbol{G})=1$) if and only if the gradients are orthogonal and equal in magnitude. Thereby, we consider the feature learning constraint $\mathcal{S}(\boldsymbol{\theta}^{\star})$ can be defined as the following optimization to eliminate instability in the training process:
$\min_{\hat{\boldsymbol{G}}}\|\boldsymbol{G}-\hat{\boldsymbol{G}}\|_{F}^{2}\quad\text{ s.t. }\hat{\boldsymbol{G}}^{\top}\hat{\boldsymbol{G}}=\boldsymbol{I}.$ (11)
@@
We can use a minimum non-zero singular value of an initial gradient matrix $\sigma(\hat{\boldsymbol{G}})=\sigma_{\min}(\boldsymbol{G})>0$ since a condition number defines the system up to scale. The problem Eq. 11 can be treated as a Procrustes problem and can be solved by performing a singular value decomposition (SVD) to $\boldsymbol{G}$ ($\boldsymbol{G}=\boldsymbol{U}\boldsymbol{\Sigma}\boldsymbol{V}^{T}$) and rescaling singular values corresponding to principal components so that they are equal to the smallest singular value:
$\hat{\boldsymbol{G}}=\sigma\boldsymbol{U}\boldsymbol{V}^{\top}=\sigma\boldsymbol{G}\boldsymbol{V}\boldsymbol{\Sigma}^{-1}\boldsymbol{V}^{T}.$ (12)
The final linear system of gradients defined by $\hat{\boldsymbol{G}}$ satisfies the optimal condition in terms of a condition number. Gradient Matrix Task-Alignment handles issues related to the stability of gradients in the two tasks addressing optimization conflicts.

### v4

@@ Abstract
Multimodal image fusion and object detection are crucial for autonomous driving. While current methods have advanced the fusion of texture details and semantic information, their complex training processes hinder broader applications. Addressing this challenge, we introduce E2E-MFD, a novel end-to-end algorithm for multimodal fusion detection. E2E-MFD streamlines the process, achieving high performance with a single training phase. It employs synchronous joint optimization across components to avoid suboptimal solutions associated to individual tasks. Furthermore, it implements a comprehensive optimization strategy in the gradient matrix for shared parameters, ensuring convergence to an optimal fusion detection configuration. Our extensive testing on multiple public datasets reveals E2E-MFD’s superior capabilities, showcasing not only visually appealing image fusion but also impressive detection outcomes, such as a 3.9% and 2.0% $\text{mAP}_{50}$ increase on horizontal object detection dataset M3FD and oriented object detection dataset DroneVehicle, respectively, compared to state-of-the-art approaches.
@@ 1 Introduction
Precise and reliable object parsing is critical in fields such as autonomous driving [1] and remote sensing monitoring [2]. Relying solely on visible sensors can lead to inaccuracies in object recognition in challenging environments, like inclement weather conditions. Visible-infrared image fusion [3; 4; 5; 6] as a typical common multimodal fusion (MF) task addresses these challenges by leveraging complementary information from different modalities, leading to the rapid development of various multimodal image fusion techniques [7; 8; 9; 10; 11]. Techniques like CDDFuse [12] and DIDFuse [13] employ a two-step process where a MF network is trained initially, followed by training an object detection (OD) network with the results from the MF network to assess fusion effectiveness separately. Although deep neural networks have significantly enhanced the ability to learn representations across modalities, resulting in promising multimodal fusion outcomes, the focus has predominantly been on producing visually appealing images. This emphasis often overlooks the improvement of downstream high-level visual tasks, such as enhanced object parsing, which remains a substantial hurdle.
@@
Recent studies have devoted into designing joint learning methods that integrate fusion networks with high-level tasks such as object detection [14] and segmentation [15; 16]. The synergy between MF and OD in Multimodal Fusion Detection (MFD) methods has emerged as a vibrant area of research. This partnership allows MF to produce richer, more informative images, enhancing OD performance, while OD contributes valuable object semantic insights to MF, aiming to accurately locate and identify objects in a scene. Typically, MFD networks adopt a cascaded design where joint optimization techniques [17] use the OD network to guide the MF network toward creating images that facilitate easier object detection. Notably, Zhao et. al [18] introduced a joint learning method for multimodal fusion detection, incorporating meta-feature embedding from OD to improve fusion by generating semantic object features through meta-learning simulation. Despite these advancements, as highlighted in Figure 1, significant challenges persist: 1) Current optimization approaches rely on a multi-step, progressively joint method, compromising efficiency; 2) These methods overly focus on leveraging OD information for fusion enhancement, leading to difficulty in parameter balancing and susceptibility to local optima of individual tasks. Therefore, the quest for a unified feature set that simultaneously caters to each task remains formidable.
@@
In this paper, we introduce E2E-MFD, an end-to-end algorithm for multimodal fusion detection, designed to seamlessly blend detailed image fusion and object detection from coarse to fine levels. E2E-MFD facilitates the interaction of intrinsic features from both domains through synchronous joint optimization, allowing for a streamlined, one-stage process. To reconcile fine-grained details with semantic information, we propose the novel concept of an Object-Region-Pixel Phylogenetic Tree (ORPPT) coupled with a coarse-to-fine diffusion processing (CFDP) mechanism. This approach is inspired by the natural process of visual perception, tailored to meet the specific needs of MF and OD. Furthermore, we introduce a Gradient Matrix Task-Alignment (GMTA) technique to fine-tune the optimization of shared components, thereby minimizing the adverse impacts traditionally associated with inherent optimization challenges. This ensures an efficient convergence towards an optimal set of fusion detection weights, enhancing both the accuracy and efficacy of multimodal fusion detection.
@@
Our contributions in this paper are highlighted as follows: (1) We present E2E-MFD, a pioneering approach to efficient synchronous joint learning, innovatively integrating image fusion and object detection into a single-stage, end-to-end framework. This methodology significantly enhances the outcomes of both tasks. (2) We introduce a novel GMTA technique, designed to evaluate and quantify the impacts of the image fusion and object detection tasks. This aids in optimizing the training process’s stability and ensures convergence to an optimal configuration of fusion detection weights. (3) Through comprehensive experimentation on image fusion and object detection, we demonstrate the efficacy and robustness of our proposed method.
@@ 2.1 Multimodal Fusion Object Detection
Due to the powerful nonlinear fitting capabilities of deep neural networks, deep learning has made significant progress in low-level vision tasks, particularly in image fusion tasks [19; 20; 7; 21; 22; 23; 24]. Early efforts [9; 25; 13; 26; 27; 28] tended to achieve excellent fusion results by adjusting network structures or loss functions, overlooking the fact that image fusion should aim to improve the performance of downstream application tasks. Fusion images with good quality metrics may be suitable for human visual perception but may not be conducive to practical application tasks [16; 29]. Some research has acknowledged this issue, Yuan et al[30] utilized various aligned modalities for improved oriented object detection [31; 32; 33; 34] to address the challenge of cross-modal weakly misalignment in aerial visible-infrared images. Liu et al. [17] proposing a joint learning method, pioneered the exploration of the MF and OD combination methods. Then the optimization of the loss function for segmentation [16] and detection [14] has been validated to be effective in guiding the generation of fused images. They consider the downstream task network as an additional constraint to assist the MF network in generating fusion results with clearer objects. Zhao et al. [18] leveraged semantic information from OD features to aid MF and perform meta-feature embedding to generate meta-features from OD features, which are then used to guide the MF network in learning pixel-level semantic information. Liu et al. [15] proposed multi-interactive feature learning architecture for image fusion and segmentation by enhancing fine-grained mapping of all the vital information between two tasks, so that the modality or semantic features can be fully mutual-interactive. However, OD considers the semantic understanding of objects, while MF and segmentation primarily focus on the pixel-level relationship between image pairs. The optimization coupling of detection and fusion tasks becomes more challenging to investigate these complementary differences, from which both fusion and detection can benefit. It is worth mentioning that these visible-infrared multimodal fusion detection methods are usually designed in a cascaded structure with tedious training steps. Researchers lack the adoption of end-to-end architectures, which would enable one-step network inference to generate credible fused images and detection results through a set of network parameters.
@@ 2.2 Multi-task Learning
Multi-task learning (MTL) [35; 36] involves the simultaneous learning of multiple tasks through parameter sharing. Prior approaches involve manually crafting the architecture, wherein the bottom layers of a model are shared across tasks [37; 38]. Some approaches tailor the architecture based on task affinity [39], while others utilize techniques such as Neural Architecture Search [40; 37; 41] or routing networks [42] to autonomously discern sharing patterns and determine the architecture. An alternative method typically combines task-specific objectives into a weighted sum [43; 44; 45]. In addition, most approaches (e.g. [46; 47; 48; 49; 50]) aim to mitigate the effects of conflicting or dominating gradients. The approach of explicit gradient modulation [48; 49; 50; 51], has demonstrated superior performance in resolving conflicts between task gradients by substituting conflicting gradients with modified, non-conflicting gradients. Inspired by the above multimodal learning methods, we introduce a Gradient Matrix Task-Alignment method to align the orthogonal components contained in image fusion and object detection tasks thereby effectively eliminating the inherent optimization barrier that exists between two tasks.
@@ 3.1 Problem Formulation
MF-OD task concentrates on generating an image that benefits emphasizing objects with superior visual perception capability. The goal of OD is to find the location and identify the class of each object in an image which can naturally provide rich semantic information along with object location information. Therefore, the motivation of OD-aware MF is to construct a novel infrared and visible image fusion framework that can benefit from the semantic information and object location information contained in OD. For this purpose, we suppose a pair of visible image $\boldsymbol{x}\in\mathbb{R}^{H\times W\times C_{x}}$ and infrared image $\boldsymbol{y}\in\mathbb{R}^{H\times W\times C_{y}}$. The optimization model is formulated as:
@@
where $\boldsymbol{t}$ represents the output of the different task network $\mathcal{N}$ with the learnable parameters $\boldsymbol{\theta}_{\mathrm{t}}$. $\mathcal{L}(\cdot)$ is a loss function to optimize the network. Previous approaches solely design image fusion or object detection networks in a cascaded way, which can only achieve outstanding results for one task. To produce visually appealing fused images alongside accurate object detection results, we jointly integrate the two tasks into a unified goal synchronously which can be rewritten as:
@@ 3.2 Architecture
where $b_{l,m}(\boldsymbol{o_{l}})$ denotes the $m^{\mathrm{th}}$ region representation and $(i,j)$ denotes the spatial location. These region-level representations are further concatenated to form the observation $B_{l}(\boldsymbol{o_{l}})=\left[b_{l,1}(\boldsymbol{o_{l}}),b_{l,2}(\boldsymbol{o_{l}}),\ldots,b_{l,M_{l}}(\boldsymbol{o_{l}})\right]$ of branch $l$. These multi-grained attentions concentrate operation on the spatial location information and the extent of the regions which are similar to the task requirements. The $B_{1}$, $B_{2}$, …, $B_{L}$ are up-sampled to keep the consistent spatial size with the pixel-level fusion features $B_{0}$. Then, the region-level fusion features $B_{1},B_{2},...,B_{L}$ are assembled by addition operation followed by a Convolution with $1\times 1$ kernel + ReLU to reduce the channel numbers. Highlighted region-level features are extracted by a Convolution with $1\times 1$ kernel + Sigmoid and then injected into pixel-level features by multiplication and addition. Finally, five Convolutions with $3\times 3$ kernel + ReLU layers are constructed to reconstruct the fusion result $u$.
@@ 3.3 Loss Function
The total loss is combined with an image fusion loss function $\mathcal{L}_{\mathrm{f}}$ and object detection loss $\mathcal{L}_{\mathrm{d}}$. $\mathcal{L}_{\mathrm{f}}$ consists of three types of losses, i.e., structure loss $\mathcal{L}_{\text{SSIM}}$, pixel loss $\mathcal{L}_{\text{pixel }}$ and gradient loss $\mathcal{L}_{\text{grad }}$. For one fused image, it should preserve overall structures and maintain a similar intensity distribution from source images. To this end, the structural similarity index (SSIM) is introduced in function:
@@
$\mathcal{L}_{\text{pixel}}^{\text{o}}=\left\|I_{m}\circ\left(\boldsymbol{u}-max(\boldsymbol{x},\boldsymbol{y})\right)\right\|_{1},\mathcal{L}_{\text{pixel}}^{\text{b}}=\left\|(1-I_{m})\circ\left(\boldsymbol{u}-mean(\boldsymbol{x},\boldsymbol{y})\right)\right\|_{1},$ (8)
where $\|\cdot\|_{1}$ stands for the $l_{1}$-norm. The operator $\circ$ denotes the elementwise multiplication, max($\cdot$) denotes the element-wise maximization, and mean($\cdot$) denotes the element-wise average operation. Therefore, the object-aware pixel loss $\mathcal{L}_{\text{pixel }}$ is defined as:
@@ 3.4 Gradient Matrix Task-Alignment
The MF and OD tasks have distinct optimization objectives. MF primarily emphasizes capturing the pixel-level relationship between image pairs, while OD incorporates object semantics within the broader context of diverse scenes. An inherent optimization barrier exists between these two tasks. We observe that the prevailing challenges in multi-task learning are arguably task dominance and conflicting gradients. We introduce a Gradient Matrix Task-Alignment (GMTA) by presenting the condition number to mitigate the undesired effects of the optimization barrier in task-shared parameters $\boldsymbol{\theta}^{\star}$ which are supposed to be balanced between the MF and OD tasks. The individual task gradients of MF and OD task are calculated by $\boldsymbol{g}_{u}=\nabla_{{\theta}^{\star}}\mathcal{L}_{u}$ and $\boldsymbol{g}_{d}=\nabla_{{\theta}^{\star}}\mathcal{L}_{d}$ in the training optimization process. The gradient matrix can be defined as $\boldsymbol{G}=\left\{\boldsymbol{g}_{u},\boldsymbol{g}_{d}\right\}$. In multi-task optimization, a cumulative gradient $\boldsymbol{g}=\boldsymbol{G}\boldsymbol{w}$ is a linear combination of task gradients and the stability of a linear system is measured by the condition number of its matrix in numerical analysis. Hence the stability of the gradient matrix is equal to the ratio of the maximum and minimum singular values (non-negative) of the corresponding matrix: $\kappa(\boldsymbol{G})=\frac{\sigma_{\max}}{\sigma_{\min}}$. Learning from the Aligned-MTL [51], the condition number is optimal ($\kappa(\boldsymbol{G})=1$) if and only if the gradients are orthogonal and equal in magnitude which means that the system of gradients has no dominance or conflicts:
$\kappa(\boldsymbol{G})=1\iff<\boldsymbol{g}_{u},\boldsymbol{g}_{d}>=1.$ (11)
The final linear system of gradients defined by $\hat{\boldsymbol{G}}$ satisfies the optimal condition in terms of a condition number. Thereby, we consider the feature learning constraint $\mathcal{S}(\boldsymbol{\theta}^{\star})$ can be defined as the following optimization to eliminate instability in the training process:
$\min_{\hat{\boldsymbol{G}}}\|\boldsymbol{G}-\hat{\boldsymbol{G}}\|_{F}^{2}\quad\text{ s.t. }\kappa(\hat{\boldsymbol{G}})=1\iff\min_{\hat{\boldsymbol{G}}}\|\boldsymbol{G}-\hat{\boldsymbol{G}}\|_{F}^{2}\quad\text{ s.t. }\hat{\boldsymbol{G}}^{\top}\hat{\boldsymbol{G}}=\boldsymbol{I}.$ (12)
@@
The problem can be treated as a Procrustes problem and can be solved by performing a singular value decomposition (SVD) to $\boldsymbol{G}$ ($\boldsymbol{G}=\boldsymbol{U}\boldsymbol{\Sigma}\boldsymbol{V}^{T}$) and rescaling singular values corresponding to principal components so that they are equal to the smallest singular value:
$\hat{\boldsymbol{G}}=\sigma\boldsymbol{U}\boldsymbol{V}^{\top}=\sigma\boldsymbol{G}\boldsymbol{V}\boldsymbol{\Sigma}^{-1}\boldsymbol{V}^{T},$ (13)
where,
$(\boldsymbol{V},\boldsymbol{\lambda})=eigh(\boldsymbol{G}^{\top}\boldsymbol{G}),$ (14)
$\boldsymbol{\Sigma}^{-1}=diag(\sqrt{1/\lambda_{max}},\sqrt{1/\lambda_{min}}),$ (15)
$eigh$ represents a function for finding eigenvectors $\boldsymbol{V}$ and eigenvalues $\boldsymbol{\lambda}$ and $diag$ stands for diagonal matrix. $\lambda_{max}$ and $\lambda_{min}$ are maximum eigenvalues and minimum eigenvalues from $\boldsymbol{\lambda}$.The stability criterion, a condition number, defines a linear system to an arbitrary position scale. To alleviate this ambiguity, we choose the largest scale that guarantees convergence to the optimum: this is a minimal singular value of an initial gradient matrix:
$\sigma=\sigma_{\min}(\boldsymbol{G})=\sqrt{\lambda_{min}}.$ (16)